OpenStudy (anonymous):

Hyp of Right Triangle is 10, 1 leg is 6, find other leg and area ?

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \bf a^2+b^2=c^2 }\) that also gives, \(\Large\color{blue}{ \bf \sqrt{c^2-b^2}=a }\)

OpenStudy (anonymous):

Okk. So

OpenStudy (solomonzelman):

So can you find the missing leg ? Or you have no clue about what I said ?

OpenStudy (anonymous):

100= 36 + b squared

OpenStudy (anonymous):

64=bsqaured

OpenStudy (solomonzelman):

also, a and b are height and base of the right triangle. Go for it :)

OpenStudy (anonymous):

b=8?

OpenStudy (solomonzelman):

yes b=8

OpenStudy (solomonzelman):

You tell me the area.

OpenStudy (anonymous):

Umm.

OpenStudy (solomonzelman):

a and b are height and base

OpenStudy (anonymous):

Which is the height ?

OpenStudy (solomonzelman):

doesn't matter

OpenStudy (anonymous):

Oh okay . So 8 x 6 =

OpenStudy (anonymous):

48 / 2 = 24

OpenStudy (anonymous):

in squared ?

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \bf A∆= \frac{b \times h}{2} }\)

OpenStudy (solomonzelman):

yes 24

OpenStudy (anonymous):

Okayy Thanks (:

OpenStudy (solomonzelman):

yw